OpenStudy (anonymous):

Which coordinates of a point is exactly 5 away from the point (7, 7) and forms a diagonal segment with that point?

OpenStudy (anonymous):

I fell stupid but i just can't get this question no matter how many times someone go's over

OpenStudy (anonymous):

^

OpenStudy (dumbcow):

(3,4) using distance formula \[d = \sqrt{(7-3)^{2}+(7-4)^{2}} = \sqrt{4^{2}+3^{2}} = \sqrt{16+9} = \sqrt{25} = 5\]

OpenStudy (anonymous):

(12, 7) (7, 12) (10, 11) (12, 12) No (3,4)

OpenStudy (dumbcow):

haha, oh well really there are an infinite number of points 5 away from (7,7) plug in each given point into distance formula and see if its 5

OpenStudy (anonymous):

ok

OpenStudy (dumbcow):

it can't be the first 2 answers (12,7) or (7,12) because they would form vertical or horizontal lines instead of a diagonal line

OpenStudy (anonymous):

So (7,12)

OpenStudy (anonymous):

ops you posted above my bad